# Bounds on detection of Bell correlations with entangled ultra-cold atoms in optical lattices under occupation defects

### v1

@@ III.2 Two-sites Bell correlations
It can be shown Müller-Rigat et al. (2021) that for any input data $\vec{M}$ and $\tilde{C}$ compatible with a LV theory (1), any $2\times 2$ positive semi-definite matrix $A$ and any $2\times 1$ vector $\vec{h}$, the following Bell inequality holds
@@
with $\vec{r}=(r_{0},r_{1})^{T}$ being the vector of all possible pairs of outcomes corresponding to the two settings $\alpha=0$ and $\alpha=1$ for a single subsystem. The positive semi-definite matrix $A$ and the vector $\vec{h}$ that minimize $L(A,h)$ can be found using a data-driven method as in Ref. Müller-Rigat et al. (2021). We find that the Bell inequality   [^1] Alternatively, the Bell inequality (22) can be represented using the $C$ matrix, namely $L=C_{00}+C_{11}-C_{01}-C_{10}-(M_{0}-M_{1})^{2}-M_{0}-M_{1}+M\geq 0$.
@@
first established in Ref. Müller-Rigat et al. (2021) for spin-squeezed states in the case of only two possible measurement outcomes, is optimal also in our case with the three possible outcomes (3) and under occupation defects.
For the case of two-sites Bell correlations, we choose $\alpha=0$ and $\alpha=1$ corresponding to the measurement, at each site $j$, of the spin components Müller-Rigat et al. (2021)
@@
To account for the occupation defects, we evaluate $M_{\alpha}$ and $\tilde{C}_{\alpha\beta}$ to determine the form of the Bell inequality using the toy model density matrix (6) under the two scenarios investigated in the physical systems presented in Secs. V and VI when $N\leq M$ and $N=M$, respectively.
@@ III.2.1 Non-unit filling
The first scenario, relevant to the system described in Sec. V, assumes that $N=pM\leq M$. Under this condition, $p<1$ indicates a non-unit filling of the lattice whose sites are empty or singly occupied. The density matrix $\hat{\rho}^{\perp}_{j}$ in (8), in this scenario, becomes
@@
for an optimal $\theta$ such that $\cos\theta_{\rm opt}=\frac{N-1}{N-p}\frac{\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}}{N-4(\Delta\hat{S}_{\vec{m}})^{2}_{\rm SS}}$ [^2] Similar to the previous case, this solution should verify $|\cos(\theta_{\rm opt})|\leq 1$, which introduces the restriction $p\leq N-(N-1)|\langle\hat{S}_{\vec{n}}\rangle\big{/}[N-4(\Delta\hat{S}_{\vec{m}})^{2}]|$. Here, $L^{\rm(V)}$ can be written in terms of the squeezing parameter $\xi^{2}=N(\Delta\hat{S}_{\vec{m}})^{2}_{\rm SS}\big{/}\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}^{2}$ and the normalized mean spin $v=\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}\big{/}(N/2)$ as $\frac{L^{\rm(V)}}{M}=1-\frac{p(N-1)}{N-p}\frac{v^{2}/4}{1-\xi^{2}v^{2}}-\frac{p(N-p)}{N-1}(1-\xi^{2}v^{2}).$ (32) . The equation (III.2.1) is represented in Fig. 2 as a function of time of the OAT dynamics for a given $N$ and for different values of the occupation probability $p$. This reveals that for a given $N$ and at each moment of time $\tau$, there is a critical value of $p$ below which, the two-site Bell correlations present in the system cannot be detected. This critical probability $p=p_{\rm c}$ is the value for which, the following $3^{\rm rd}$ order equation holds
@@
Figure 2: Variation of the non-locality witness $L_{\rm opt}/M$, given by (39) in the upper panel and (III.2.1) in the lower panel, during the one-axis twisting dynamics with $N=10^{3}$ for different values of $p$ (probability of single occupation of a given site). The flat behaviour at short times happens because of the optimization condition.
III.2.2 Doubly occupied sites
In this scenario, relevant to the system described in Sec. VI, we assume that $N=M$. Under this condition, $p<1$ indicates a redistribution of the $N$ atoms among the $M$ lattice sites resulting in the emergence of both empty and doubly occupied sites. The filling factor is one, $f=1$. Thus, the density matrix $\hat{\rho}^{\perp}_{j}$ in (8) can be written as
@@ III.2.2 Doubly occupied sites
for an optimal $\theta$ such that $\cos\theta_{\rm opt}=\frac{1}{p}\frac{\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}}{N-4(\Delta\hat{S}_{\vec{m}})^{2}_{\rm SS}}$ [^3] This solution should verify $|\cos(\theta_{\rm opt})|\leq 1$ which introduces the restriction $p\geq|\langle\hat{S}_{\vec{n}}\rangle\big{/}[N-4(\Delta\hat{S}_{\vec{m}})^{2}]|$., [^4] An alternative formula of $L^{\rm(VI)}$ can be obtained by introducing the spin squeezing parameter $\xi^{2}$ and the normalized mean spin $v$ as $\frac{L^{\rm(VI)}}{M}=1-\frac{v^{2}/4}{1-\xi^{2}v^{2}}-p^{2}(1-\xi^{2}v^{2}).$ (40) . The equation (39) is represented in Fig. 2-(a) as a function of time $\tau$ of the OAT dynamics for a given $N$ and for different values of $p$. We note that for a given $N$ and at each moment of time, there is a critical value of $p$ below which, the two-site Bell correlations present in the system cannot be detected, from (39) we obtain
@@
and the corresponding critical probability $p_{\rm c}$ approaches $p_{\rm c}=\sqrt{3}/2$. A lower value of $p$ can be obtained with a Bell inequality including onsite and two-site correlations up to the fourth-order. In the limit of large $N$, this leads to a critical occupation probability of $p_{\rm c}=1/2$ Baamara (2023).
@@ IV Application with ultra-cold atoms in optical lattices
We test the predictions of our toy model using a system consisting of $N$ ultra-cold bosonic atoms confined in an optical lattice. We focus on rubidium-87 atoms occupying two internal states, labelled $a$ and $b$, and loaded into an optical lattice potential, akin to recent experiments employing quantum gas microscopes Rubio-Abadal et al. (2019); Wei et al. (2022). The optical lattice comprises $M$ lattice sites and is considered to be in one dimension.
The system, in the lowest energy band, is conveniently considered in the Wannier functions basis Bloch et al. (2008). In the tight-binding limit, where the lattice potential exceeds the recoil energy $E_{R}=\hbar^{2}k^{2}/(2m)$, the system can be conveniently described by the Bose-Hubbard model,
@@
We explore two protocols for the dynamical generation of Bell correlations within this system as detailed in Sections V and VI, both resulting in Mott entangled states where atoms exhibit spin entanglement and are evenly distributed across the lattice. In both cases, the initial state $\hat{\rho}_{a}$ with all atoms in the internal state $|a\rangle$ is turned into a coherent superposition of $|a\rangle$ and $|b\rangle$ by applying a $\pi/2-$pulse, which is equivalent to
@@
In the first protocol, Bell correlations are generated directly in the strongly interacting Mott regime $U_{\sigma\sigma^{\prime}}\gg t$ where individual spins interact via spin-exchange interactions between neighbouring spins (46). These interactions allow for the generation of entanglement within the system either through weak anisotropy or by coupling with an inhomogeneous field. The mechanism for generating entanglement was explained and demonstrated in prior works such as Hernández Yanes et al. (2022, 2023, 2024). In the second protocol, Bell correlations are induced by atom-atom interactions in the superfluid regime $t\gg U_{\sigma\sigma^{\prime}}$ with all-to-all individual spin connections. Subsequently, these correlations are transferred to the Mott phase by an adiabatic increase of the optical lattice depth, a process described in detail in previous studies Kajtoch et al. (2018a); Płodzień et al. (2020a, 2022).
@@ V Entanglement generation in the Mott phase
is the Heisenberg XXZ model with the spin-exchange interaction parameter $J=4t^{2}/U_{ab}$ and the anisotropy parameter $\Delta=U_{ab}/U_{aa}+U_{ab}/U_{bb}-1$ Altman et al. (2003). When $\Delta=1$ the Hamiltonian takes the form of the isotropic Heisenberg XXX model and it is a natural case for rubidium-87 where $U_{aa}\approx U_{bb}\approx U_{ab}$. The anisotropy parameter $\Delta$ can be tuned by changing the values of interaction strengths using either Feshbach resonances or by shifting optical lattice potentials for states $a$ and $b$. The collective spin operators are just a summation over the individual ones, $\hat{S}_{\sigma}=\sum_{j=1}^{M}\hat{s}^{(j)}_{\sigma}$ for $\sigma=x,y,z,\pm$. The tunnelling term in (45) is relevant whenever the filling factor $f=N/M$ is not one, meaning there are holes (empty sites) in the system and $N\leq M$.
Let us consider the case of zero temperature when the initial state $\hat{\rho}_{a}$ is in the Mott regime. In the ideal case, we have $\hat{\rho}_{a}=\bigotimes_{j=1}^{M}|a\rangle\langle a|_{j}$, indicating that at each lattice site there is an atom in the internal state $|a\rangle$. However, in the presence of holes, the on-site state can be $|0\rangle\langle 0|_{j}$ if it is not occupied. In the given experimental realisation, the number of holes can be arbitrary as well as their positions. In our simulations, we evolve single realizations with initially each site being empty or singly occupied, and then average over many realizations to evaluate expectation values. Specifically, we generate a random number $x_{j}\in(0,1]$ for each lattice site, obtaining
@@
where $\bar{\theta}(x)$ is the Heaviside step function. The density matrix $\hat{\rho}_{a}$ describing $N_{r}$ realizations is obtained through a simple averaging,
$\hat{\rho}_{a}=\frac{1}{N_{r}}\sum_{r=1}^{N_{r}}\left[\bigotimes_{j=1}^{M}|\Psi_{x}\rangle\langle\Psi_{x}|_{j}\right],$ (48)
over the set of on-site random numbers $r=\{x_{j}^{1},x_{j}^{2},\cdots,x_{j}^{{N_{r}}}\}$ for all $j\in\{1,...M\}$. One can show, that in the continuous limit, we obtain
$\hat{\rho}_{a}=\bigotimes_{j=1}^{M}\left(f|a\rangle\langle a|_{j}+(1-f)|0\rangle\langle 0|_{j}\right),$ (49)
given that all the sites are equivalent. Therefore, the filling factor $f$ can be identified with the parameter $p$ of the external state of the toy model (6) provided that $\hat{\rho}_{j}^{\perp}=|0\rangle\langle 0|_{j}$. Owing to the presence of holes, the average number of atoms $N$ differs from the number of lattice sites $M$ when $f\neq 1$, namely $N=fM$.
@@
$\begin{split}\hat{\rho}_{a}&=f^{M}\bigotimes_{j=1}^{M}|a\rangle\langle a|_{j}\\ &+f^{M}\left(\frac{1-f}{f}\right)^{1}\left(|0\rangle\langle 0|_{1}\bigotimes_{j=1}^{M-1}|a\rangle\langle a|_{j}+\mathcal{P}_{1}\right)\\ &+f^{M}\left(\frac{1-f}{f}\right)^{2}\left(|0\rangle\langle 0|_{1}\otimes|0\rangle\langle 0|_{2}\bigotimes_{j=1}^{M-2}|a\rangle\langle a|_{j}+\mathcal{P}_{2}\right)\\ &+\cdots,\end{split}$ (50)
@@
$\hat{\rho}_{a}=f^{M}\left[\bigotimes_{j=1}^{M}|a\rangle\langle a|_{j}+\sum_{\sigma=1}^{M}\left(\frac{1-f}{f}\right)^{\sigma}\Sigma_{\sigma}\left[\hat{\rho}_{\sigma}\right]\right],$ (51)
@@
and $\hat{\rho}_{\rm ini}^{(0)}=e^{-i\hat{S}_{y}\pi/2}\left(\bigotimes_{j=1}^{M}|a\rangle\langle a|_{j}\right)e^{i\hat{S}_{y}\pi/2}$. Therefore, any expectation values of the operator $\hat{O}$ can be evaluated as

### v2

@@ III.2 Two-sites Bell correlations
It can be shown [29] that for any input data $\vec{M}$ and $\tilde{C}$ compatible with a LV theory (1), any $2\times 2$ positive semi-definite matrix $A$ and any $2\times 1$ vector $\vec{h}$, the following Bell inequality holds
@@
with $\vec{r}=(r_{0},r_{1})^{T}$ being the vector of all possible pairs of outcomes corresponding to the two settings $\alpha=0$ and $\alpha=1$ for a single subsystem. The positive semi-definite matrix $A$ and the vector $\vec{h}$ that minimize $L(A,h)$ can be found using a data-driven method as in Ref. [29]. We find that the Bell inequality   [^1] Alternatively, the Bell inequality (22) can be represented using the $C$ matrix in (19), namely $L=C_{00}+C_{11}-C_{01}-C_{10}-(M_{0}-M_{1})^{2}-M_{0}-M_{1}+M\geq 0$.
@@
first established in Ref. [29] for spin-squeezed states in the case of only two possible measurement outcomes, is optimal also in our case with the three possible outcomes (3) and under occupation defects.
For the case of two-sites Bell correlations, we choose $\alpha=0$ and $\alpha=1$ corresponding to the measurement, at each site $j$, of the spin components [29]
@@
We now evaluate $M_{\alpha}$ and $\tilde{C}_{\alpha\beta}$ using the toy model density matrix (6), for the physical situation presented in Sec. V where $N\leq M$, and that in Sec.VI where $N=M$.
@@ III.2.1 Non-unit filling
The first scenario, relevant to the system described in Sec. V, assumes that $N=pM\leq M$ where $N$ is the number of atoms and $M$ is the number of sites. Under this condition, $p<1$ indicates a non-unit filling of the lattice whose sites are empty or singly occupied. The density matrix $\hat{\rho}^{\perp}_{j}$ in (8), in this scenario, becomes
@@
where the subscript ${\rm SS}$ refers to an expectation calculated in the $\hat{\rho}_{\rm SS}$ state (7).
@@
for an optimal $\theta$ such that $\cos\theta_{\rm opt}=\frac{N-1}{N-p}\frac{\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}}{N-4(\Delta\hat{S}_{\vec{m}})^{2}_{\rm SS}}$ [^2] This solution should verify $|\cos(\theta_{\rm opt})|\leq 1$, which introduces the restriction $p\leq N-(N-1)|\langle\hat{S}_{\vec{n}}\rangle\big{/}[N-4(\Delta\hat{S}_{\vec{m}})^{2}]|$. Here, $L^{\rm(V)}$ can be written in terms of the squeezing parameter $\xi^{2}=N(\Delta\hat{S}_{\vec{m}})^{2}_{\rm SS}\big{/}\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}^{2}$ and the normalized mean spin $v=\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}\big{/}(N/2)$ as $\frac{L^{\rm(V)}}{M}=1-\frac{p(N-1)}{N-p}\frac{v^{2}/4}{1-\xi^{2}v^{2}}-\frac{p(N-p)}{N-1}(1-\xi^{2}v^{2}).$ (32) . The equation (III.2.1) is represented in Fig. 2 (upper panel) as a function of time of the OAT dynamics for a given $N$ and for different values of the occupation probability $p$. This reveals that for a given $N$ and at each moment $\tau$, there is a critical value of $p$ below which, the two-site Bell correlations present in the system cannot be detected. This critical probability $p=p_{\rm c}$ is the value for which, the following $3^{\rm rd}$ order equation holds
@@
Figure 2: Variation of the non-locality witness $L_{\rm opt}/M$, given by (III.2.1) in the upper panel and (39) in the lower panel, during the one-axis twisting dynamics with $N=10^{3}$ for different values of $p$ (probability of single occupation of a given site). The flat behaviour at short times is a consequence of the optimization condition.
III.2.2 Double occupancies and empty sites
In a second scenario, relevant to the system described in Sec. VI, we assume that $N=M$ where $N$ is the number of atoms and $M$ is the number of sites. Under this condition, $p<1$ indicates a redistribution of the $N$ atoms among the $M$ lattice sites resulting in the emergence of both empty and doubly occupied sites. The filling factor is one, $f=1$. Thus, the density matrix $\hat{\rho}^{\perp}_{j}$ in (8) can be written as
@@ III.2.2 Double occupancies and empty sites
for an optimal $\theta$ such that $\cos\theta_{\rm opt}=\frac{1}{p}\frac{\langle\hat{S}_{\vec{n}}\rangle_{\rm SS}}{N-4(\Delta\hat{S}_{\vec{m}})^{2}_{\rm SS}}$ [^3] This solution should verify $|\cos(\theta_{\rm opt})|\leq 1$ which introduces the restriction $p\geq|\langle\hat{S}_{\vec{n}}\rangle\big{/}[N-4(\Delta\hat{S}_{\vec{m}})^{2}]|$., [^4] An alternative formula of $L^{\rm(VI)}$ can be obtained by introducing the spin squeezing parameter $\xi^{2}$ and the normalized mean spin $v$ as $\frac{L^{\rm(VI)}}{M}=1-\frac{v^{2}/4}{1-\xi^{2}v^{2}}-p^{2}(1-\xi^{2}v^{2}).$ (40) . The equation (39) is represented in Fig. 2 (lower panel) as a function of time $\tau$ of the OAT dynamics for a given $N$ and for different values of $p$. We note that for a given $N$ and at each moment, there is a critical value of $p$ below which, the two-site Bell correlations present in the system cannot be detected, from (39) we obtain
@@
and the corresponding critical probability $p_{\rm c}$ approaches $p_{\rm c}=\sqrt{3}/2$. A lower value of $p$ can be obtained with a Bell inequality including onsite and two-site correlations up to the fourth order. In the limit of large $N$, this leads to a critical occupation probability of $p_{\rm c}=1/2$ [47].
@@ IV Application with ultra-cold atoms in optical lattices
We test the predictions of our toy model using a system consisting of $N$ ultra-cold bosonic atoms confined in an optical lattice. We focus on rubidium-87 atoms occupying two internal states, labelled $a$ and $b$, and loaded into an optical lattice potential, akin to recent experiments employing quantum gas microscopes [48, 49]. The optical lattice comprises $M$ lattice sites and is considered to be in one dimension.
Using the Wannier functions basis [50], when the system is in the lowest energy band and in the tight-binding limit, the lattice potential exceeding the recoil energy $E_{R}=\hbar^{2}k^{2}/(2m)$, the dynamics is conveniently described by the Bose-Hubbard Hamiltonian
@@
We explore two protocols for the dynamical generation of Bell correlations within this system as detailed in Sections V and VI, both resulting in Mott entangled states where atoms exhibit spin entanglement and are distributed across the lattice. In both cases, the initial state $\hat{\rho}_{a}$ with all atoms in the internal state $|a\rangle$ is turned into a coherent superposition of $|a\rangle$ and $|b\rangle$ by applying a $\pi/2-$pulse, which is equivalent to
@@
In the first protocol, Bell correlations are generated directly in the strongly interacting Mott regime $U_{\sigma\sigma^{\prime}}\gg t$ where individual spins interact via spin-exchange interactions between neighbouring spins (46). These interactions allow for the generation of entanglement within the system either through weak anisotropy or by coupling with an inhomogeneous field. The mechanism for generating entanglement was explained and demonstrated in prior works such as [42, 51, 41]. In the second protocol, Bell correlations are induced by atom-atom interactions in the superfluid regime $t\gg U_{\sigma\sigma^{\prime}}$ with all-to-all individual spin connections and transferred to the Mott phase by an adiabatic increase of the optical lattice depth, a process described in detail in previous studies [31, 32, 28].
@@ V Entanglement generation in the Mott phase
is the Heisenberg XXZ model with the spin-exchange interaction parameter $J=4t^{2}/U_{ab}$ and the anisotropy parameter $\Delta=U_{ab}/U_{aa}+U_{ab}/U_{bb}-1$ [52]. When $\Delta=1$ the Hamiltonian takes the form of the isotropic Heisenberg XXX model and it is a natural case for rubidium-87 where $U_{aa}\approx U_{bb}\approx U_{ab}$. The anisotropy parameter $\Delta$ can be tuned by changing the values of interaction strengths using either Feshbach resonances or by shifting optical lattice potentials for states $a$ and $b$. The collective spin operators are just a summation over the individual ones, $\hat{S}_{\sigma}=\sum_{j=1}^{M}\hat{s}^{(j)}_{\sigma}$ for $\sigma=x,y,z,\pm$. The tunnelling term in (45) is relevant whenever the filling factor $f=N/M$ is not one, meaning there are holes (empty sites) in the system and $N\leq M$.
Let us consider the case of zero temperature when the initial state $\hat{\rho}_{a}$ is in the Mott regime. In the ideal case, we have $\hat{\rho}_{a}=\bigotimes_{j=1}^{M}|a\rangle\langle a|_{j}$, indicating that at each lattice site there is an atom in the internal state $|a\rangle$. However, in the presence of holes, the on-site state can be $|0\rangle\langle 0|_{j}$ if it is not occupied. In a given experimental realisation, the number of holes can be arbitrary as well as their positions. In our simulations, we evolve single realizations with initially each site being empty or singly occupied, and then average over many realizations to evaluate expectation values. Specifically, we generate a random number $x_{j}\in(0,1]$ for each lattice site, obtaining
@@
where $f$ is the filling factor and $\bar{\theta}(x)$ is the Heaviside step function. The density matrix $\hat{\rho}_{a}$ describing $N_{r}$ realizations is obtained through a simple averaging,
$\hat{\rho}_{a}=\frac{1}{N_{r}}\sum_{l=1}^{N_{r}}\left[\bigotimes_{j=1}^{M}|\Psi_{x}\rangle_{j}{}_{j}\langle\Psi_{x}|\right],$ (48)
over the set of random numbers $x=\{x_{1}^{l},x_{2}^{l},\cdots,x_{M}^{{l}}\}$ with $l\in\{1,...N_{r}\}$. One can show, that in the continuous limit, we obtain
$\hat{\rho}_{a}=\bigotimes_{j=1}^{M}\left(f|a\rangle_{j}{}_{j}\langle a|+(1-f)|0\rangle_{j}{}_{j}\langle 0|\right),$ (49)
given that all the sites are equivalent. Therefore, the filling factor $f$ can be identified with the parameter $p$ of the external state of the toy model (6) provided that $\hat{\rho}_{j}^{\perp}=|0\rangle_{j}{}_{j}\langle 0|$. Owing to the presence of holes, the average number of atoms $N$ differs from the number of lattice sites $M$ when $f\neq 1$, namely $N=fM$.
@@
$\begin{split}\hat{\rho}_{a}&=f^{M}\bigotimes_{j=1}^{M}|a\rangle_{j}{}_{j}\langle a|\\ &+f^{M}\left(\frac{1-f}{f}\right)^{1}\left(|0\rangle_{1}{}_{1}\langle 0|\bigotimes_{j=1}^{M-1}|a\rangle_{j}{}_{j}\langle a|+\mathcal{P}_{1}\right)\\ &+f^{M}\left(\frac{1-f}{f}\right)^{2}\left(|0\rangle_{1}{}_{1}\langle 0|\otimes|0\rangle_{2}{}_{2}\langle 0|\bigotimes_{j=1}^{M-2}|a\rangle_{j}{}_{j}\langle a|+\mathcal{P}_{2}\right)\\ &+\cdots,\end{split}$ (50)
@@
$\hat{\rho}_{a}=f^{M}\left[\bigotimes_{j=1}^{M}|a\rangle_{j}{}_{j}\langle a|+\sum_{\sigma=1}^{M}\left(\frac{1-f}{f}\right)^{\sigma}\Sigma_{\sigma}\left[\hat{\rho}_{\sigma}\right]\right],$ (51)
@@
and $\hat{\rho}_{\rm ini}^{(0)}=e^{-i\hat{S}_{y}\pi/2}\left(\bigotimes_{j=1}^{M}|a\rangle_{j}{}_{j}\langle a|\right)e^{i\hat{S}_{y}\pi/2}$. Therefore, any expectation values of the operator $\hat{O}$ can be evaluated as